OpenStudy (anonymous):

Calvin must prove that this equation is an identity. tanx=(1/sinxcosx)-(cosx/sinx)

OpenStudy (anonymous):



OpenStudy (anonymous):

1/(sinxcosx) - cosx/sinx = 1/(sinxcosx) - cos^2x/(sinxcosx) = (1-cos^2x)/(sinxcosx) = sin^2x/(sinxcosx) = sinx/cosx = tanx

OpenStudy (anonymous):

*where cos^2x + sin^2x = 1 (from trig identities)

OpenStudy (anonymous):

\[\tan x = \sin x/\cos x = \sin ^{2}x/(\sin x \cos x) = (1-\cos ^{2}x)/(\cos x \sin x)\] separating the numerator intor into 2 terms u'll get right hand side part